The 1995 Nobel Prize in Chemistry was shared by Paul Crutzen, F. Sherwood Rowland, and Mario Molina for their work concerning the formation and decomposition of ozone in the stratosphere. Rowland and Molina hypothesized that chlorofluorocarbons (CFCs) in the stratosphere break down upon exposure to UV radiation, producing chlorine atoms. Chlorine was previously identified as a catalyst in the breakdown of ozone into oxygen gas.

Using the enthalpy of reaction for two reactions with ozone, determine the enthalpy of reaction for the reaction of chlorine with ozone.

(1)ClO(g)+O3(g)⟶Cl(g)+2O2(g)Δ?∘rxn=−122.8 kJ

(2)2O3(g)⟶3O2(g)Δ?∘rxn=−285.3 kJ

(3)O3(g)+Cl(g)⟶ClO(g)+O2(g)   Δ?∘rxn= ?

Δ?∘rxn=______kJ
Answer #1

Lets number the reaction as 1, 2, 3 from top to bottom
required reaction should be written in terms of other reaction
This is Hess Law
required reaction can be written as:
reaction 3 = -1 * (reaction 1) +1 * (reaction 2)

So, ΔHo rxn for required reaction will be:
ΔHo rxn = -1 * ΔHo rxn(reaction 1) +1 * ΔHo rxn(reaction 2)
= -1 * (-122.8) +1 * (-285.3)
= -162.5 KJ
Answer: -162.5 KJ